# Cross-Subject DD: A Cross-Subject Brain-Computer Interface Algorithm

### v1

@@ Abstract
Brain-computer interface (BCI) based on motor imagery (MI) enables direct control of external devices by decoding the electroencephalogram (EEG) generated in the brain during imagined movements. However, due to inter-individual variability in brain activity, existing BCI models exhibit poor adaptability across subjects, thereby limiting their generalizability and widespread application. To address this issue, this paper proposes a cross-subject BCI algorithm named Cross-Subject DD (CSDD), which constructs a universal BCI model by extracting common features across subjects. The specific methods include: 1) training personalized models for each subject; 2) transforming personalized models into relation spectrums; 3) identifying common features through statistical analysis; and 4) constructing a cross-subject universal model based on common features. The experiments utilized the BCIC IV 2a dataset, involving nine subjects. Eight of these subjects were selected for training and extracing the common features, and the cross-subject decoding performance of the model was validated on the remaining subject. The results demonstrate that, compared with existing similar methods, our approach achieves a 3.28% improvement in performance. This paper introduces for the first time a novel method for extracting pure common features and constructing a universal cross-subject BCI model, thereby facilitating broader applications of BCI technology.
*Correspondence: gangliu_@zzu.edu.cn
@@ 1 Introduction
Brain-computer interface (BCI) enables the direct translation of brain signals into control commands for computers or external devices, offering vast potential across diverse fields such as neural rehabilitation, neuroscience, augmented reality, and more[1][2][3]. Central to the functionality of BCI technology is the real-time decoding of neural activity, often relying on techniques like electroencephalography (EEG), functional magnetic resonance imaging (fMRI), and other brain imaging modalities[4]. Among these technologies, EEG-based BCI has emerged as a leading area of research and application, primarily due to its non-invasive nature, low cost, and strong real-time capabilities[5]. Owing to the enhancement of computational capabilities and continuous advancements in machine learning technologies, BCI systems have witnessed a significant improvement in performance, leading to breakthroughs in various practical applications.
However, despite these advancements, the widespread deployment of BCI technology in real-world scenarios continues to face numerous challenges. One of the foremost obstacles is the issue of poor cross-subject generalization[6]. Current BCI models are typically trained for individual users and often struggle to generalize to new users. This is primarily due to inherent differences in brain structures, neural activity patterns, and electrophysiological signals across individuals[7]. As a result, extensive user-specific adaptation and retraining are required, necessitating large datasets and considerable computational resources. These limitations significantly impede the commercialization and broader adoption of BCI technology.
To address this challenge, research on cross-subject BCI has gained increasing attention. An ideal cross-subject BCI system should be capable of operating seamlessly across multiple users without the need for extensive individual calibration. However, developing an effective cross-subject model remains a significant challenge due to the substantial inter-subject variability in brain electrophysiological activity. Existing methods often struggle to extract stable and distinctive common features across individuals, leading to poor adaptability and limited performance in cross-subject[8].
Several approaches have been proposed to improve cross-subject BCI performance, including deep learning (DL), transfer learning (TL), domain adaption (DA), data augmentation, among others. Among these, deep learning has been particularly prominent, due to its powerful feature extraction capabilities[9][10][11]. Models like EEGNet, introduced by Lawhern et al., leverage deep learning to automatically learn hierarchical representations from raw, noisy, and non-stationary EEG signals, thus enhancing generalization across subjects while maintaining interpretability even with limited data[9]. In addition to deep learning, transfer learning has been widely explored as a way to reduce the data collection burden for new users. This method adapts pre-trained models from source subjects to new individuals, facilitating more efficient adaptation[12][13]. Zhang et al. proposed HDNN-TL, an end-to-end hybrid deep neural network combining convolutional neural networks (CNN) and long short-term memory (LSTM) with transfer learning[14]. The network was fine-tuned using rectified linear units (RELU) in the fully connected layer to accommodate new tasks. The testing showed that the framework achieved satisfactory results for new subjects, reducing both time and training data requirements in the motor imagery task. Domain adaptation is a key technique in transfer learning, with the primary goal of reducing the data distribution gap between the source domain and the target domain. This enables the model to maintain good generalization performance, even when the target domain has little or no labeled samples. Fu et al. developed the SCDAN model[15], utilizing domain adaptation techniques to learn shared neural activation patterns, thereby improving the transferability of motor imagery EEG decoding. To enhance model robustness, data augmentation techniques have been used to generate synthetic EEG samples that mimic inter-subject variability[16][17].
Despite these advancements, several challenges remain. While deep learning models are effective at identifying common features across subjects using large-scale annotated data, they are still susceptible to overfitting due to the variability inherent in individual subjects[18]. Transfer learning and domain adaptation can enhance model performance by incorporating subject-specific features, but they introduce additional computational complexity and data requirements[19][20]. Furthermore, although data augmentation techniques can simulate inter-subject variability, the synthetic samples may lack physiological realism, which limits their overall effectiveness[21].
A major limitation of most existing approaches lies in focusing on subject-specific adaptation, which fine-tunes models for individual users rather than systematically extracting stable and distinguishable neural representations across subjects. We argue that the most direct approach to cross-subject adaptation is to extra common features among individuals and construct a universal model to enable broader applicability. However, to date, there is no well-established method for systematically extracting inter-subject commonalities in BCI applications, which significantly limits the generalization ability of current models.
To address these issues, this paper introduces a novel approach that explicitly distinguishes between subject-specific and common features. We propose a cross-subject BCI algorithm based on spectral analysis, referred to as Cross-Subject DD (CSDD). Unlike conventional methods, our core innovation lies in extracting common components across subjects to construct a widely applicable BCI model. Our framework consists of four key steps. (see Fig. 1)
(1) Training personalized BCI models for each subject.
(2) Transforming personalized models into relation spectrums.
(3) Applying statistical analysis to identify common features present across multiple subjects.
(4) Constructing a generalized BCI model based on the extracted common features.
Through this approach, our model enhances adaptability and decoding accuracy for new users without requiring extensive subject-specific training data.
The primary contributions of this paper are as follows.
(1) This paper proposes for the first time a method for extracting both common and subject-specific features in human intentions, along with visualization and selection techniques.
(2) This paper proposes for the first time a model based on common features across subjects during human intention generation, demonstrating that common feature modeling enhances model generalization across subjects.
(3) Filter: This paper proposes a novel method for decomposing a system into relation spectrums and then “extracting” the necessary components to reconstruct a new system (system filter). This method is similar to the Fourier filter in signal analysis, which decomposes a complex signal into its frequency spectrum, and from which the desired components are extracted to reconstruct the signal.
@@
The remainder of this paper is organized as follows. Section 2 provides a detailed description of the proposed CSDD model. Section 3 introduces the datasets and experimental setups. Section 4 presents the experimental results. Section 5 discusses the implications of our findings. Section 6 concludes the paper.
Figure 1: Flow chart of the proposed cross-subject decoding model(CSDD).
@@
Figure 2: Overall model architecture of the proposed CSDD for cross-subject decoding. The proposed CSDD model consists of four major steps. 1. SSTL-PF. 2. TRM-RS. 3. ECF-SA. 4. BCSDM-CF. Ns: the number of subjects. m: the number of common relation items. O: the number of outputs. Q: query. K: key. V: value.
2 CSDD
CSDD is a model designed to extract common features and build models based on these features. The model aims to solve the cross-subject generalization problem by learning common features across subjects while filtering out subject-specific features. The framework of the CSDD model consists of four key components, which are explained in detail in the following sections: subject-specific transfer learning based on pre-training and fine-tuning (SSTL-PF), transformation of personalized models based on relation spectrums (TPM-RS), extraction of common features based on statistical analysis (ECF-SA), building cross-subject decoding model based on common features (BCSDM-CF). The overall flowchart of the CSDD model is illustrated in Fig. 1.
2.1 Subject-specific transfer learning based on pre-training and fine-tuning (SSTL-PF)
In recent years, transfer learning has become a widely used approach to address the problem of generalizing brain-computer interface models across different subjects. One of the key techniques within transfer learning is pretraining on a large, diverse dataset followed by fine-tuning on a subject-specific dataset. This approach has been shown to be effective in extracting common features across subjects while adapting to the individual characteristics of each subject. At this stage, subject-specific transfer learning (SSTL-PF) is employed, which includes two sub-components: universal feature extraction (UFE) and dendrite net filter (DD Filter).
2.1.1 Universal feature extraction (UFE) as shown in Fig. 2(a)
The Universal Feature Extraction (UFE) process is based on an end-to-end convolutional neural network architecture, adapted from the model proposed by Ma et al.[22] , with modifications to support subsequent fine-tuning. During the pre-training phase, EEG data from $N_{\mathrm{s}}-1$ subjects are used to train the universal model.
Raw EEG signals from $N_{\mathrm{s}}-1$ subjects are represented as $E\in\mathbb{R}^{N\times C\times T}$, where $N$ denotes the number of trials, $C$ represents the number of electrodes, and $T$ indicates the number of sample points. These signals are then processed through a series of temporal convolutional layers with kernel sizes of (1, 15), (1, 25), (1, 51), and (1, 65) to extract features at multiple temporal scales. The resulting feature maps are then passed through spatial convolutional layers with a kernel size of (C, 1), where C represents the number of electrode channels, to capture spatial relationships across the different EEG channels. Following the spatial convolution, the feature maps undergo dimensionality reduction and feature selection via varpool and avgpool layers, which help retain important information while reducing computational complexity. The pooled features are then reshaped and input into a shared self-attention module, which utilizes a multi-head attention mechanism and feed-forward networks to weight and integrate features, enhancing their representational capacity. Next, the features are processed by a convolutional encoder using a (1, 2) kernel, followed by batch normalization and activation (ELU) to prepare the model for the final classification task. The processed features are then passed through a fully connected (FC) layer to generate the final output.
To facilitate the subsequent fine-tuning with DD Filter, a linear dimensionality reduction layer is added before the FC layer, reducing the feature space to 32 dimensions and performing normalization. Let $X\in\mathbb{R}^{32\times 1}$ represent the resulting 32-dimensional feature vector, which serves as the input for the subsequent processing stages.
2.1.2 DD Filter as shown in Fig. 2(b)
Dendrite Net, or DD, is a fundamental machine learning algorithm and a white-box model used for classification, regression, and system identification[23][24]. It is characterized by its white-box nature, controllable accuracy, and low computational complexity, which enhance its generalization performance. It resembles the Taylor series expansion and can converge to a function (or system) that closely approximates the target. Unlike traditional neural networks, it does not require non-linear functions but instead utilizes multiple DD modules to construct logical expressions among input features.
In this study, the author team has made significant innovations to the previously proposed DD model, introducing for the first time a filtering framework based on the DD algorithm, namely the DD Filter. The innovation of the DD Filter lies in its ability to transform the trained DD model into a relation spectrum. By utilizing statistical analysis, the DD Filter effectively identifies and retains common relation items across subjects while filtering out subject-specific features, which are treated as noise. This section focuses on the model representation of the DD Filter, with additional details on its implementation provided in Sections 2.2 and 2.3.
Within the CSDD framework, the DD Filter plays a crucial role in converting extracted EEG features into meaningful representations, facilitating subsequent filtering processes. During the fine-tuning stage, all layers of the pretrained model, except the fully connected layer, are frozen. The fully connected layer is replaced with the DD module for feature learning, and the model is subsequently trained using data from the current subject. This results in $N_{\mathrm{s}}-1$ fine-tuned models, also known as personalized models. It receives a 32-dimensional feature vector from the preceding layer, obtained through a universal feature extraction process.
@@ 2.1 Subject-specific transfer learning based on pre-training and fine-tuning (SSTL-PF)
The DD Filter in this paper consists of three key components: the dimensionality reduction module, the DD module, and the linear module. To enhance computational efficiency while preserving essential information, a dimensionality reduction module projects this feature vector onto a 16-dimensional space. The formula of the dimensionality reduction module is as follows.
$G_{1}=W_{1}^{T}X$ (1)
@@
where $G_{1}\in\mathbb{R}^{16\times 1}$ expresses the outputs of the dimensionality reduction module. $W_{1}\in\mathbb{R}^{32\times 16}$ is the weight matrix. $X\in\mathbb{R}^{32\times 1}$ expresses the inputs of the dimensionality reduction module and the outputs of the UFE module.
DD modules aim at no-linear mapping. The degree of interaction items is adjusted by the number of DD modules. Each DD module is equivalent to a first-order approximation, and stacking multiple DD modules results in a higher-order approximation. Therefore, by adjusting the number of DD modules, the mapping capability of DD can be controlled, preventing overfitting or underfitting. The DD module in this paper is expressed as follows.
$G_{2}=W_{G_{1}}G_{1}\circ G_{1}$ (2)
@@
where $G_{1}\in\mathbb{R}^{16\times 1}$ and $G_{2}\in\mathbb{R}^{16\times 1}$ express the inputs and outputs of the DD module, respectively. $W_{G_{1}}\in\mathbb{R}^{16\times 16}$ is the weight matrix. $\circ$ expresses Hadamard product.
The linear module, whose dimensions are determined by the number of output categories. The formula is expressed as follows.
$G^{\text{SSTL-PFoutputs}}=W_{2}^{T}G_{2}$ (3)
@@
where $G_{2}\in\mathbb{R}^{16\times 1}$ express the outputs of the DD module. $W_{2}\in\mathbb{R}^{16\times O}$ is the weight matrix. $G^{SSTL-PFoutputs}\in\mathbb{R}^{O\times 1}$ is the outputs of SSTL-PF. $O$ is the number of categories determines.
2.2 Transformation of personalized models based on relation spectrums (TPM-RS)
The relation spectrum is a polynomial expression based on the dendrite network[23][25][26]. Unlike traditional spectral analysis (such as Fourier transform), it emphasizes the relation items between the input signals and the output decoding. By analyzing the polynomial and its coefficients, the relation spectrum provides a deeper insight into neural decoding. In this study, the fine-tuned personalized models are transformed into relation spectrums through symbolic simplification. This transformation allows us to analyze feature interactions and identify cross-subject stable components for building generalized BCI decoding module.
2.2.1 Symbolic simplification of the DD model

### v2

# System Filter-Based Common Components Modeling for Cross-Subject EEG Decoding

@@ Abstract
Brain–computer interface (BCI) technology enables direct communication between the brain and external devices through electroencephalography (EEG) signals. However, existing decoding models often mix common and personalized components, leading to interference from individual variability that limits cross-subject decoding performance. To address this issue, this paper proposes a system filter that extends the concept of signal filtering to the system level. The method expands a system into its spectral representation, selectively removes unnecessary components, and reconstructs the system from the retained target components, thereby achieving explicit system-level decomposition and filtering. We further integrate the system filter into a Cross-Subject Decoding framework based on the System Filter (CSD-SF) and evaluate it on the four-class motor imagery (MI) task of the BCIC IV 2a dataset. Personalized models are transformed into relation spectrums, and statistical testing across subjects is used to remove personalized components. The remaining stable relations, representing common components across subjects, are then used to construct a common model for cross-subject decoding. Experimental results show an average improvement of 3.28% in decoding accuracy over baseline methods, demonstrating that the proposed system filter effectively isolates stable common components and enhances model robustness and generalizability in cross-subject EEG decoding.
{IEEEkeywords}
Brain-computer interface (BCI), Electroencephalogram (EEG), Cross-subject decoding, System filter, Motor imagery (MI)
Figure 1: The research goals of this paper. (a) Previous decoding models mix common and personalized components, leading to poor generalization for new subjects. (b) The proposed approach filters out personalized components and builds a pure cross-subject model based only on common components, enabling stable decoding.
Figure 2: Conceptual analogy between the proposed system filter and signal filtering. (a) Signal filtering. A complex signal with noise is mapped into the frequency domain via Fourier transform and decomposed into multiple components. Filtering out unnecessary frequencies (e.g., 120 Hz, 180 Hz) yields a clean waveform containing only the target components (e.g., 20 Hz, 50 Hz). (b) System filter. A complex system, represented as a sum of polynomial terms, can be transformed into a spectral representation. By removing unnecessary terms (e.g., $x_{1}$, $x_{3}$, $x_{i}^{2}$) and retaining only the target ones, a purer and more interpretable system is obtained.
@@ 1 Introduction
[…]
Brain-computer interface (BCI) systems decode electroencephalography (EEG) signals to directly control external devices, bypassing muscular and neural pathways. They have been widely applied in neurorehabilitation, motor assistance, and intelligent device control [1, 2, 3, 4]. However, cross-subject BCI decoding remains a significant challenge, as inter-subject variability substantially reduces the generalization capability of BCI models[5, 6, 7]. This variability, which arises from subject-specific differences in brain activity patterns, makes it difficult for models to generalize effectively across different individuals[8]. As a result, improving the robustness of BCI models and enhancing their generalization across subjects is a critical issue that needs to be addressed.
Conventional BCI decoding models are hybrid systems, combining two components: common components, which capture patterns shared across subjects, and personalized components, which capture subject-specific variations [9, 10, 11]. During training, these two elements are implicitly mixed into a single model. As illustrated in Fig. 1(a), the purple region denotes common components, while the blue and green regions represent personalized components corresponding to different subjects. When a new subject is introduced, their EEG signals inevitably contain new personalized components, which were not present during training. These unseen components interfere with the trained hybrid model, leading to mismatches and reducing decoding accuracy [12, 13].
To address this problem, transfer learning (TL) has become the dominant approach[14, 15, 16, 17]. In a typical TL pipeline, a model is first pre-trained on data from multiple subjects and then fine-tuned with a small amount of data from a new subject. Fine-tuning emphasizes personalized features, adapting the model to the individual. However, this approach does not eliminate the interference caused by subject-specific components. While it improves within-subject performance, it does not solve the problem of cross-subject generalization, as the underlying hybrid structure remains sensitive to subject-specific variations.
This paper proposes a new approach that focuses solely on common components to construct a universal cross-subject decoding model, as shown in Fig. 1(b). Rather than adapting a model to individual subjects by fine-tuning, the proposed method aims to filter out personalized components and retain only the common components shared across subjects. This approach seeks to eliminate the interference of subject-specific variability and improve the robustness and generalization of BCI decoding across different subjects.
To achieve this goal, we propose a system filter, inspired by the principles of signal filtering, as illustrated in Fig. 2. For example, Fourier transform maps a time-domain signal into the frequency domain, allowing for the selective removal of unnecessary frequency bands while retaining the essential signal[18]. Similarly, the system filter transforms a complex system into a polynomial representation, maps it into spectral space, and selectively removes unnecessary components, finally reconstructs the system from the retained components. The human brain can be viewed as a biological system [19], where EEG recordings reflect brain activity that translates into specific intentions or motor commands. This EEG-based system serves as a biological information transmission channel, mapping neural activity to intended actions[20]. By applying the system filter, we aim to isolate the stable, common patterns of brain activity across subjects, removing personalized variations. These retained common components are then used to construct a universal model, enabling more accurate and generalized cross-subject decoding.
The main contributions of this paper are as follows.
1 This paper proposes a system filter that transforms the system into spectrum representation, filters out unnecessary components, and retains only the target components, yielding a system composed solely of target components (see Fig. 2). This method parallels signal filtering, where signals are decomposed in the frequency domain to remove noise.
2 This is the first study to apply the system filter to cross-subject motor imagery electroencephalography (MI-EEG) decoding tasks. By separating common components from subject-specific ones and constructing a universal model, experimental results demonstrate that the proposed system filter improves generalization, enhancing cross-subject decoding performance.
@@
The remainder of this paper is organized as follows. Section 2 provides a detailed description of the system filter and its application in cross-subject MI-EEG decoding. Section 3 introduces experimental setups. Section 4 presents the experimental results. Section 5 discusses the implications of our findings. Section 6 concludes the paper.
2 Methods
2.1 System Filter
The proposed system filter follows a three-step procedure: expansion, filtering, and reconstruction, which are described in detail below.
2.1.1 Expansion
Any finite-dimensional system can be represented in the form of a polynomial expansion. Mathematically, based on the completeness of the polynomial basis, all internal relations of the system can be mapped onto a set of polynomial terms. Let $f:\mathbb{R}^{d}\to\mathbb{R}$ denote a system with input vector $x=(x_{1},x_{2},\ldots,x_{d})\in\mathbb{R}^{d}$. To explicitly reveal its internal structure, the system is expanded into a relation spectrum.
$f(x)=\sum_{k=1}^{K}c_{k}\,\phi_{k}(x),$ (1)
where $\phi_{k}(x)$ denotes the $k$-th polynomial basis function, and $c_{k}$ is the corresponding coefficient. This expansion provides a complete and interpretable decomposition of the system into separable relation items.
2.1.2 Filtering
In the relation spectrum space, some relation items are necessary to the target output, while others correspond to unnecessary components. To filter out the unnecessary components, a selection operator $\mathcal{S}$ is defined as follows.
$\mathcal{S}(f)=\sum_{k\in\mathcal{I}}c_{k}\,\phi_{k}(x)$ (2)
where $\mathcal{I}\subseteq\{1,2,\ldots,K\}$ denotes the index set of retained items. This operation performs “filtering” in the spectral space, yielding a purified representation of the target components.
2.1.3 Reconstruction
After filtering, the system is reconstructed by recombining the retained relation items.
$f^{*}(x)=\mathcal{S}(f)(x)$ (3)
This reconstruction yields a system composed solely of target components, eliminating the interference of unnecessary terms while achieving system-level decomposition and filtering in a compact, interpretable, and mathematically principled manner.
@@ 2.1.3 Reconstruction
Figure 3: Overall model architecture of the proposed CSD-SF for cross-subject EEG decoding for MI. The proposed CSD-SF consists of four major modules. 1. APM. 2. ERS. 3. ECC. 4. CCM. m: the number of common relation items. O: the number of outputs. Q: query. K: key. V: value.
2.2 Application of the System Filter to Cross-Subject MI-EEG Decoding
To validate the effectiveness of the system filter in cross-subject EEG decoding tasks, this section applies it to MI-EEG decoding tasks, referred to as CSD-SF. The method is mainly divided into four steps. (1) Acquisition of the personalized model (APM), (2) Expansion of the personalized model (EPM), (3) Extraction of common components (ECC), (4) Construction of a cross-subject decoding model based on common components (CCM) (see Fig. 3).
2.2.1 Dataset
The BCIC IV 2a dataset contains EEG signals from 22 electrodes, recorded from nine healthy subjects over two sessions conducted on two different days. Each session consists of 288 four-second motor imagery trials per subject, involving the imagined movements of the left hand, right hand, feet, and tongue. This study uses data from 0 to 4 seconds. Before release, the signals were sampled at 250 Hz and bandpass filtered between 0.5 Hz and 100 Hz. The original dataset uses 288 trials from the first session (“T”) for training and 288 trials from the second session (“E”) for testing. In the cross-subject scenario, the original dataset needs to be re-split by subject using a leave-one-subject-out approach for modeling based on common components. Consequently, nine datasets (A01-A09) are obtained, each containing 576 trials (288 trials $\times$ 2 sessions) for testing one subject, with the remaining 4608 trials (288 trials $\times$ 2 sessions $\times$ 8 subjects) used for training. This procedure yields nine folds, ensuring that each subject serves as the unseen test subject once. The BCIC IV 2a dataset is publicly available as described in [21]. The experimental paradigm of the BCIC IV 2a dataset is shown in Fig. 4.
Figure 4: The experimental paradigm for BCIC IV 2a
2.2.2 Acquisition of the Personalized Model (APM)
To obtain individualized models, hybrid features extracted by the hybrid feature extraction (HFE) framework were fed into the dendrite network (DD), a model proposed by our research group[29, 30, 31], for training. As a white-box model with a linear–Hadamard structure, DD naturally corresponds to a finite-order polynomial expansion, providing an explicit representation that enables subsequent transformation into relation spectrums for filtering.
HFE: The HFE module is implemented using an end-to-end convolutional neural network adapted from Ma et al.[28] . EEG data from eight subjects are used to train a universal model. Raw EEG signals are denoted as $E\in\mathbb{R}^{N\times C\times T}$, where $N$ is the number of trials, $C$ is the number of electrodes, and $T$ indicates the number of sample points. Temporal convolutional layers with kernel sizes $(1,15),(1,25),(1,51),(1,65)$ extract multi-scale temporal features, which are subsequently processed by spatial convolution across channels $(C,1)$ to capture inter-channel dependencies. The outputs are further refined by varpool and avgpool layers, and integrated using a multi-head self-attention mechanism. Finally, a 32-dimensional normalized feature vector is obtained via linear dimensionality reduction and used as the input to the DD model.
$\boldsymbol{X}\in\mathbb{R}^{32\times 1}$ (4)
@@ 2.2.2 Acquisition of the Personalized Model (APM)
DD: First, a dimensionality reduction module projects the input vector $\boldsymbol{X}\in\mathbb{R}^{32\times 1}$ into a 16-dimensional subspace to enhance computational efficiency while preserving essential information.
$\boldsymbol{G}_{1}=\boldsymbol{W}_{1}^{T}\boldsymbol{X}$ (5)
@@
where $\boldsymbol{G}_{1}\in\mathbb{R}^{16\times 1}$ expresses the outputs of the dimensionality reduction module. $\boldsymbol{W}_{1}\in\mathbb{R}^{32\times 16}$ is the weight matrix. $\boldsymbol{X}\in\mathbb{R}^{32\times 1}$ expresses the inputs of the dimensionality reduction module and the outputs of the UFE module. The DD module in this paper is expressed as follows.
$\boldsymbol{G}_{2}=\boldsymbol{W}_{G_{1}}\boldsymbol{G}_{1}\circ\boldsymbol{G}_{1}$ (6)
@@
where $\boldsymbol{G}_{1}\in\mathbb{R}^{16\times 1}$ and $\boldsymbol{G}_{2}\in\mathbb{R}^{16\times 1}$ express the inputs and outputs of the DD module, respectively. $\boldsymbol{W}_{G_{1}}\in\mathbb{R}^{16\times 16}$ is the weight matrix. $\circ$ expresses Hadamard product. Finaly, the linear module, whose dimensions are determined by the number of output categories.
$\boldsymbol{G}^{\text{APM\_outputs}}=\boldsymbol{W}_{2}^{T}\boldsymbol{G}_{2}$ (7)
@@
where $\boldsymbol{G}_{2}\in\mathbb{R}^{16\times 1}$ express the outputs of the DD module. $\boldsymbol{W}_{2}\in\mathbb{R}^{16\times O}$ is the weight matrix. $\boldsymbol{G}^{\text{APM\_outputs}}\in\mathbb{R}^{O\times 1}$ is the outputs of APM. $O$ is the number of categories determines.
2.2.3 Expansion of the Personalized Model (EPM)-Expansion